OpenStudy (nuccioreggie):

If (44)x = 432, what is the value of x? 4 7 8 28

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (nuccioreggie):

I think its 8

OpenStudy (welshfella):

divide 432 by 44

OpenStudy (solomonzelman):

is that \(\large\color{black}{ \displaystyle (44)^x=432 }\) or \(\large\color{black}{ \displaystyle (44)\times x=432 }\)

OpenStudy (solomonzelman):

and no the answer isn't a whole numb.

OpenStudy (solomonzelman):

the quickest step for (44)^x is take log base 44, and use a calculator for the left side. for (44) times x, just divide as welshfella has told you too

OpenStudy (solomonzelman):

to*

Nnesha (nnesha):

it's \[\huge\rm (4^4)^x =4^{32}\] right ?

OpenStudy (nuccioreggie):

So A

Nnesha (nnesha):

which one is it ? is this ur question ??\[\huge\rm (4^4)^x =4^{32}\]

OpenStudy (nuccioreggie):

Yea

Nnesha (nnesha):

alright so now YOU tell me the answer remember we did this yesterday ?

OpenStudy (nuccioreggie):

I think its B

Nnesha (nnesha):

why ???

OpenStudy (solomonzelman):

Nnesha, that is the question ?

OpenStudy (nuccioreggie):

Yea it is

Nnesha (nnesha):

yeah i guess^

Nnesha (nnesha):

what exponent rule you should apply ?

OpenStudy (nuccioreggie):

Ok can u do wat u did yester day wen u asked me what do i multiply to get this number,so it will be much eisier

Nnesha (nnesha):

you already know ^^^ it's a same method but with different numbers

OpenStudy (solomonzelman):

ok... next time use an equation editor or denote an exponent using a caret, the ^ (Shift+6)

Nnesha (nnesha):

\[\huge\rm x^m \times x^n = x^{m \times n}\] m=4 n=x m times n = 32 substitute 4 and then solve for n

OpenStudy (nuccioreggie):

Ok can u do wat u did yester day wen u asked me what do i multiply to get this number,so it will be much eisier

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @nuccioreggie Ok can u do wat u did yester day wen u asked me what do i multiply to get this number,so it will be much eisier \(\color{blue}{\text{End of Quote}}\) \(\color{blue}{\text{Originally Posted by}}\) @Nnesha \[\huge\rm x^m \times x^n = x^{m \times n}\] m=4 n=x m times n = 32 substitute 4 and then solve for n \(\color{blue}{\text{End of Quote}}\)

OpenStudy (nuccioreggie):

8

Nnesha (nnesha):

check ur answer \[\huge\rm (4^4)^8 = 4^{ 4 \times 8} = 4^{?}\]

OpenStudy (nuccioreggie):

8 ?

OpenStudy (nuccioreggie):

Im thinking

OpenStudy (solomonzelman):

u don't accept messages, so I will say this right here. you maid a mistake when you quoted the poster. The law of multiplying x^n times x^m gives x^(m+n)

OpenStudy (solomonzelman):

\(\Large\color{black}{ \displaystyle \rm x^m\times x^n=x^{m+n} }\)

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @SolomonZelman u don't accept messages, so I will say this right here. you maid a mistake when you quoted the poster. The law of multiplying x^n times x^m gives x^(m+n) \(\color{blue}{\text{End of Quote}}\) i posted the wrong one facepalm >.<

Nnesha (nnesha):

\[\huge\rm (x^m)^n = x^{m \times n}\]

OpenStudy (nuccioreggie):

Im so confused the answer is B right??

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \[\huge\rm x^m \times x^n = x^{m \times n}\] m=4 n=x m times n = 32 substitute 4 and then solve for n \(\color{blue}{\text{End of Quote}}\) i mean to post this exponent rule \[\huge\rm (x^m)^n = x^{m \times n}\] \[(4^4)^x =4^32 \]

Nnesha (nnesha):

4^{32}**

Nnesha (nnesha):

check your answer \[\huge\rm (4^4)^7 = 4^{32}\] is this right 4 times 7 =32? \[\large\rm (4^4)^8 =4^{32}\] both sides are equal ?? 4 times 8=32 ?? \[\huge\rm (4^4)^4 = 4^{32}\] 4 times 4 = 32 ?

Nnesha (nnesha):

when you multiply both exponents you should get 32 got it ? :-)

OpenStudy (nuccioreggie):

Wow thats not even an answer choice thats the wrong answer

Nnesha (nnesha):

huh?? which one ??

OpenStudy (nuccioreggie):

If (44)x = 432, what is the value of x? 4 7 8 28

OpenStudy (nuccioreggie):

That question

Nnesha (nnesha):

i'm saying if you u don't know exponent rules then you can substitute all answer choices for x to check your answer

OpenStudy (nuccioreggie):

What is the answer to this question

Nnesha (nnesha):

you already know \[\huge\rm (4^4)^x =4^{32}\] 4 times x = 32 you got it right!

Nnesha (nnesha):

4 times what =32 ?

OpenStudy (nuccioreggie):

is the answer 8

Nnesha (nnesha):

yeah i got the same answer 4 times 8 =32